OpenStudy (anonymous):

Choose the point which is a solution to the system of inequalities: y less than or equal to –2x + 3 y < one half x – 4 options: A. 4, –6) B. (–4, 6) C. (4, 6) D. (–4, –6)

OpenStudy (anonymous):

@Hero can u help me please?

hero (hero):

Test each point directly by plugging them in.

hero (hero):

For example, let's try plugging the first point (4, -6) into each inequality: \(y \le -2x + 3 \) \( -6\le -2(4) + 3 \) \(-6 \le -8 + 3\) \(-6 \le -5\) True \(y \le \frac{x}{2} - 4\) \(-6 \le \frac{4}{2} - 4\) \(-6 \le 2 - 4\) \(-6 \le -2\) True Well, I only wanted to show you an example, but the first one ended up being true for both expressions. Looks like (4, -6) is the correct point.

OpenStudy (anonymous):

lol thanks